Find to 4 decimal places the following binomial probabilities using the normal approximation. a. n-150, p0.44, P(x-73) P(x-73)- b. n-100, ?-0.56, P(52 x 59) P(52 Sx s 59)- P(x 2 40)- d. n-105, ?-0.72, P(x 78) Click if you would like to Show Work for this question: Open Show Work
Answer #1

a) X ~ Binornial(n-150, p=0.44) Binomial can be approximated to normal with: σ = v/np(1-p) = V 150 * (0.44) (1-0.44) 6.079474- 0.0330 P(X=73) = 0.0330 Using excel instead of z-table gives: 0.0338312797 b) X ~ Binomial(n = 100, p = 0.56) Binomial canXBinomial (n - 90,p- 0.1) ial can be approximated to normal with: np-90 * 0.41-36.9 μ p) = V/90 * (0.41) (1-0.41) 4.665940 4.d) XBinomial (n - 105, p0.72) ial can be approximated to normal with: μ-np-105 * 0.72 = 75.6 σ = V np(1-p-V 105 * (0.72) (1-0